The non-volatile, non-electrolyte TNT (trinitrotoluene), C7H5N3O6
(227.10g/mol), is soluble in benzene C6H6. Calculate the osmotic
pressure generated when 14.3g of TNT are dissolved in 242ml of a
benzene solution at 298 K.
1
Expert's answer
2020-11-02T08:51:15-0500

Solution:

The osmotic pressure of a solution is proportional to the molar concentration of the solute particles in solution.

Π = i × CM × R × T

where:

  • Π is the osmotic pressure,
  • i is the van't Hoff factor, 
  • CM is the molar concentration,
  • R is the ideal gas constant (0.0821 L atm mol-1 K-1),
  • T is the temperature in Kelvin.


Molar concentration (CM) is the number of moles of solute per liter of solution, which can be calculated using the following equation: CM = mol solute / L of solution.

CM = n / V = m / (M × V)

CM(TNT) = (14.3 g) / (227.10 g mol-1 × 0.242 L) = 0.2602 mol L-1 = 0.26 mol L-1

CM(TNT) = 0.26 mol L-1


The van't Hoff factor (i) for the solute is 1 because it is a nonelectrolyte.

Hence,

Π = CM × R × T

Π = 0.26 mol L-1 × 0.0821 L atm mol-1 K-1 × 298 K = 6.36 atm

Π = 6.36 atm


Answer: the osmotic pressure (Π) is 6.36 atm.